Can someone please help me find the equation of the line containing the given pair of points?
(1,5) and (3,6)
y=?

Answer by rfer(12644)

first find slope
m=(6-5)/(3-1)
m=1/2
-----------------
now find y intercept
y=mx+b
6=.5*3+b
6=1.5+b
b=4.5
------------
y=mx+b
y=1/2x+4.5